# Evidence of Memory Effects in the Dynamics of Two-Level System Defect Ensembles Using Broadband, Cryogenic Transient Dielectric Spectroscopy

### v2

@@ III Results
In this section, we provide a more in-depth analysis of the transient dielectric response from the sample in Fig. 3d containing photoresist on sapphire, which is shown in Fig. 5. We start by plotting the transient dielectric response in Fig. 5aii showing long-timescale oscillatory collapse and revival behavior of the emitted signal, which indicates information backflow and memory effects [59, 60]. This can be seen clearly in Fig. 5ai, which displays the response at a single frequency of GHz. To quantify the frequency components of the oscillations observed in the time traces in Fig. 5a, we perform a FFT of all the data traces to generate Fig. 5bii. Again, a particular line cut at the same frequency is shown in Fig. 5bi. The FFT shows clear V-shaped features and horizontal bands, particularly in the 3.65–4.0 GHz and 4.25–4.65 GHz ranges. These V-shapes arise from off-resonant emission by dominant eigenfrequencies of the ensembles of TLS defects, made visible by the broadband response of the waveguide [61]. Numerical simulations of a driven, interacting four-spin system (Appendix G, Fig. 13) confirm this behavior, namely the appearance of V-shaped arcs in the FFT, which sharpen with longer pulse durations, and converge toward the bare eigenmode frequencies.
@@
where is the intensity of the transmitted pulse, and plot it in Fig. 5c. provides us with a map of temporal correlations and displays evident non-exponential, oscillatory decay, which is a hallmark signature of memory effects or non-Makovian dynamics of the ensembles of TLS defects. Finally, we utilize the computed to extract using Eq. 14, which we show in Fig. 5d. Here, consistent with , we observe non-monotonic and negative , highlighted by the cyan contours at zero in Fig. 5d, exhibiting partial overlap with the aforementioned V-shapes in some regions of the FFT. This behavior indicates physics beyond linear response and further supports the presence of memory effects and re-emission in the ensembles of TLS defects [63, 64].
@@ IV Conclusions and Outlook
This work introduces BCTDS, a technique to probe TLS defects that supports fast and modular mounting of various samples, without invoking narrow-band probes like resonators or qubits. Our measured cryogenic dielectric spectroscopy experiments reveal distinct transient features and memory effects as demonstrated by the and plots in Fig. 5c,d, respectively. Such memory effects may indicate that the dynamics are non-Markovian, which may be detrimental for quantum error correction if manifested in quantum processors [65, 66, 54]. We believe that much of the memory effects arise from the interplay between coherent TLS defect interactions and driving and measuring in a broadband environment. By analyzing the frequency and temporal correlations of large ensembles of TLS defects in isolated materials, we take initial steps toward uncovering the atomistic structure of TLS defects. Next, we aim to introduce an orthogonal SMA port to rotate the polarization of the drive and perform the same dielectric spectroscopy experiments, subject to different drive strengths. By integrating over different drive polarizations, we can better understand the frequency distribution and dipole moments of the TLS defects in various samples.
@@ Appendix C Cryogenic Dielectric Spectroscopy of the Same Sample, For Different Cooldowns
Two-level system (TLS) defects are known to shift in frequency and spatial location after thermal cycling [68], due to rearrangements in the disordered environment that reshape the local energy landscape. To corroborate this signature, we perform dielectric spectroscopy experiments 15 days apart on the same sample subject to a thermal cycle to 300 K in Fig. 8. As expected, we see a significantly different driven and transient dielectric response in Fig. 8a,b, highlighted by the difference plot in Fig. 8c. In future work, we intend to perform systematic studies of the temperature dependence closer to 10 mK. Such studies may provide insight into the observed ‘universal dielectric response’, and potentially probe the transition of quantum TLS defects to the broad classical dielectric response [69, 70].
@@ Appendix D Dielectric Response in Transmission and Reflection
For ideal, lossless systems, transmission and reflection contain equivalent information as incident power must be conserved. In our experiment, where dissipation is present, reflection provides complementary information about dielectric loss during the pulsing region. We make measurements on the same samples as in Appendix C but place a cryogenic circulator at the waveguide input to measure both the transmission (Fig. 9a) and reflection (Fig. 9b) signals. In the transient region, emission from the dielectric samples is detected almost equally in the transmission and reflection channels, as the samples are positioned equidistant from the two antenna pins, and the readout circuits are identical. As a result, the collected signal shows minimal difference, as shown in Fig. 9c. A short vertical white line is added to indicate pulse bandwidth f.
@@ Appendix E Electric Field Amplitude Calibration
The waveguide design ensures that the transmitted signal is relatively constant over a bandwidth of approximately 3-6 GHz (Fig. 2b). However, the samples under test are coupled in the middle of the waveguide, which can be frequency-dependent as illustrated in Fig. 10. We leverage HFSS simulations to predict the electric field amplitude at the sample plane over our frequency bandwidth (blue dot and lines in Fig. 10b). To compensate for this variation, we apply a frequency-dependent scaling factor (red solid line in Fig. 10b). The impact of this calibration is evident when comparing the transient spectra of a sapphire sample with AlOx deposit before (Fig. 10c) and after (Fig. 10d) applying the correction. Significantly, most of the key features in the data remain. Still, such calibrations can ensure uniform electric field densities as a function of frequency and may prove critical in future experiments to deduce properties of individual ensembles of TLS defects.
@@ Appendix G Qualitative Comparison to Numerical Simulations
Explicitly, the Fourier transform of the dipole response function is expressed via the Floquet Green’s functions, , as
@@
Figure 12 shows numerical simulations of the collective excitation dynamics and Floquet quasi-energies in a two-spin system. Panel a displays the time evolution of the collective population, obtained by solving the Lindblad master equation with fixed dipole-dipole coupling, drive amplitude, pulse duration (marked by a horizontal dashed line), collective dissipation, and two distinct TLS frequencies (indicated by vertical dashed lines). During the drive, clear Rabi-Chevron patterns emerge, while following the pulse, pronounced interference and long-lived ring-down oscillations are observed. Figure 12b shows the Fourier amplitude of the full time-domain simulation as a function of drive frequency. The V-shaped features converging toward the bare TLS frequencies reflect detunings between the drive and the individual transitions. These structures highlight the coherent excitation conditions that give rise to the long-lived ring-downs observed in panel a. Figure 12c displays the first four Floquet quasi-energies computed during the driving period. The quasi-energy spectrum exhibits characteristic dips near the driving frequencies, which correspond to resonances that enhance coherent excitation trapping. These resonances manifest as long-lived ring-down dynamics in the time domain, indicating that the system temporarily stores energy in quasi-stationary Floquet states with suppressed decay, thus prolonging the coherent oscillations observed after the pulse ends.
Figure 13 presents numerical simulations of a driven, interacting four-spin network under a finite-duration pulse. In the experiment, the emitted signal is measured via homodyne detection of the in-phase (I) and quadrature (Q) components, which are proportional to the collective dipole radiation from the system. This output field corresponds to the collective lowering operator , so the measured power reflects the expectation value , directly connecting to the simulated observable . The upper panels show the time-domain evolution of this collective excitation as a function of drive frequency for increasing pulse durations, where pronounced interference patterns emerge in the ring-down following the pulse, as shown in Fig. 13a. Figure 13b displays the corresponding Fourier spectra, which reveal how the spectral content of the ring-down evolves with increasing pulse duration. For short pulses, the abrupt temporal cutoff introduces large energy uncertainty, resulting in broad sidebands. As the pulse duration increases, the drive becomes more spectrally selective, allowing coherent oscillations among near-resonant modes to build up over time. This temporal coherence leads to narrower sidebands in the Fourier spectrum, which progressively sharpen and approach the bare transition frequencies of the system. This sharpening gives rise to the observed V-shaped structure in the frequency domain that we observe experimentally.
@@
Acknowledgements.
Bert Harrop graciously diced the devices at Princeton. Simon Agnew helped perform the aluminum oxide deposition using an atomic layer deposition tool in William Scheideler’s lab at Dartmouth. Sofia Abrego and Chloe Buschmann helped review and polish the manuscript. We also acknowledge stimulating conversations with Kater Murch, Miles Blencowe, Geoffroy Hautier, Chandrasekhar Ramanathan, and Valla Fatemi. Startup funds from the Thayer School of Engineering at Dartmouth primarily supported this work. While the scientific focus of this work is quite distinct from existing funding, M.F. would like to express gratitude for the support of science within his group through the DARPA Young Faculty Award No. D23AP00192 and NSF PHY-2412555, and hopes that projects like this continue to be made possible by grants from the US federal government. The views and conclusions in this manuscript are those of the authors. They should not be interpreted as representing the official policies, expressed or implied, of DARPA, the NSF, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.Author Contributions
Competing Interests
The authors declare no competing interests.
Additional information
Correspondence and request for materials

### v3

@@ III.3 Further Investigation of Cryogenic Transient Dielectric Spectroscopy Results
In this section, we provide a more in-depth analysis of the transient dielectric response from the sample in Fig. 3(f) containing photoresist on sapphire, which is shown in Fig. 5. We start by plotting the transient dielectric response in Fig. 5(aii) showing long-timescale oscillatory collapse and revival behavior of the emitted signal, which indicates information backflow and memory effects [59, 60]. This can be seen clearly in Fig. 5(ai), which displays the response at a single frequency of GHz. To quantify the frequency components of the oscillations observed in the time traces in Fig. 5(a), we perform a FFT of all the data traces to generate Fig. 5(bii). Again, a particular line cut at the same frequency is shown in Fig. 5(bi). The FFT shows clear V-shaped features and horizontal bands, particularly in the 3.65–4.0 GHz and 4.25–4.65 GHz ranges. These V-shapes arise from off-resonant emission by dominant eigenfrequencies of the ensembles of TLS defects, made visible by the broadband response of the waveguide [61]. Numerical simulations of a driven, interacting four-spin system (Appendix G, Fig. 13) confirm this behavior, namely the appearance of V-shaped arcs in the FFT, which sharpen with longer pulse durations, and converge toward the bare eigenmode frequencies.
@@
where is the intensity of the transmitted pulse, and plot it in Fig. 5(c). provides us with a map of temporal correlations and displays evident non-exponential, oscillatory decay, which is a hallmark signature of memory effects or non-Makovian dynamics of the ensembles of TLS defects. Finally, we utilize the computed to extract using Eq. 14, which we show in Fig. 5(d). Here, consistent with , we observe non-monotonic and negative , highlighted by the cyan contours at zero in Fig. 5(d), exhibiting partial overlap with the aforementioned V-shapes in some regions of the FFT. This behavior indicates physics beyond linear response and further supports the presence of memory effects and re-emission in the ensembles of TLS defects [63, 64].
@@ IV Conclusions and Outlook
This work introduces BCTDS, a technique to probe TLS defects that supports fast and modular mounting of various samples, without invoking narrow-band probes like resonators or qubits. Our measured cryogenic dielectric spectroscopy experiments reveal distinct transient features and memory effects as demonstrated by the and plots in Fig. 5(c), 5(d), respectively. Such memory effects may indicate that the dynamics are non-Markovian, which may be detrimental for quantum error correction if manifested in quantum processors [65, 66, 54]. We believe that much of the memory effects arises from the interplay between coherent TLS defect interactions and driving and measuring in a broadband environment. By analyzing the frequency and temporal correlations of large ensembles of TLS defects in isolated materials, we take initial steps toward uncovering the atomistic structure of TLS defects. Next, we aim to introduce an orthogonal SMA port to rotate the polarization of the drive and perform the same dielectric spectroscopy experiments, subject to different drive strengths. By integrating over different drive polarizations, we can better understand the frequency distribution and dipole moments of the TLS defects in various samples.
@@
Acknowledgements.
Bert Harrop graciously diced the devices at Princeton. Sofia Abrego and Chloe Buschmann helped review and polish the manuscript. We also acknowledge stimulating conversations with Kater Murch, Miles Blencowe, Geoffroy Hautier, Chandrasekhar Ramanathan, and Valla Fatemi. Startup funds from the Thayer School of Engineering at Dartmouth primarily supported this work. While the scientific focus of this work is quite distinct from existing funding, M.F. would like to express gratitude for the support of science within his group through the DARPA Young Faculty Award No. D23AP00192 and NSF PHY-2412555, and hopes that projects like this continue to be made possible by grants from the US federal government. The views and conclusions in this manuscript are those of the authors. They should not be interpreted as representing the official policies, expressed or implied, of DARPA, the NSF, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.Author Contributions
Competing Interests
The authors declare no competing interests.
Additional information
Correspondence and request for materials
@@ Appendix C Cryogenic Dielectric Spectroscopy of the Same Sample, For Different Cooldowns
Two-level system (TLS) defects are known to shift in frequency and spatial location after thermal cycling [68], due to rearrangements in the disordered environment that reshape the local energy landscape. To corroborate this signature, we perform dielectric spectroscopy experiments 15 days apart on the same sample subject to a thermal cycle to 300 K in Fig. 8. As expected, we see a significantly different driven and transient dielectric response in Fig. 8(a) and 8(b), highlighted by the difference plot in Fig. 8(c). In future work, we intend to perform systematic studies of the temperature dependence closer to 10 mK. Such studies may provide insight into the observed ‘universal dielectric response’, and potentially probe the transition of quantum TLS defects to the broad classical dielectric response [69, 70].
@@ Appendix D Dielectric Response in Transmission and Reflection
For ideal, lossless systems, transmission and reflection contain equivalent information as incident power must be conserved. In our experiment, where dissipation is present, reflection provides complementary information about dielectric loss during the pulsing region. We make measurements on the same samples as in Appendix C but place a cryogenic circulator at the waveguide input to measure both the transmission (Fig. 9(a)) and reflection (Fig. 9(b)) signals. In the transient region, emission from the dielectric samples is detected almost equally in the transmission and reflection channels, as the samples are positioned equidistant from the two antenna pins, and the readout circuits are identical. As a result, the collected signal shows minimal difference, as shown in Fig. 9(c). A short vertical white line is added to indicate pulse bandwidth f.
@@ Appendix E Electric Field Amplitude Calibration
The waveguide design ensures that the transmitted signal is relatively constant over a bandwidth of approximately 3-6 GHz, shown in Fig. 2(b). However, the samples under test are coupled in the middle of the waveguide, which can be frequency-dependent, as illustrated in Fig. 10(a). We leverage HFSS simulations to predict the average electric field amplitude at the sample plane over our frequency bandwidth , plotted in blue dots and lines in Fig. 10(b). To compensate for this variation, we apply a frequency-dependent scaling factor (red solid line). The impact of this calibration is evident when comparing the transient spectra of a sapphire sample with an AlOx deposit before (Fig. 10(c)) and after (Fig. 10(d)) applying the correction. Significantly, most of the key features in the data remain. Still, such calibrations can ensure uniform electric field densities as a function of frequency and may prove critical in future experiments to deduce properties of individual ensembles of TLS defects.
@@ Appendix G Qualitative Comparison to Numerical Simulations
Explicitly, the Fourier transform of the dipole response function is expressed via the Floquet Green’s functions, as
@@
Figure 12 shows numerical simulations of the collective excitation dynamics and Floquet quasi-energies in a two-spin system. Panel (a) displays the time evolution of the collective population, obtained by solving the Lindblad master equation with fixed dipole-dipole coupling, drive amplitude, pulse duration (marked by a horizontal dashed line), collective dissipation, and two distinct TLS frequencies (indicated by vertical dashed lines). During the drive, clear Rabi-Chevron patterns emerge, while following the pulse, pronounced interference and long-lived ring-down oscillations are observed. Figure 12(b) shows the Fourier amplitude of the full time-domain simulation as a function of drive frequency. The V-shaped features converging toward the bare TLS frequencies reflect detunings between the drive and the individual transitions. These structures highlight the coherent excitation conditions that give rise to the long-lived ring-downs observed in panel (a). Figure 12(c) displays the first four Floquet quasi-energies computed during the driving period. The quasi-energy spectrum exhibits characteristic dips near the driving frequencies, which correspond to resonances that enhance coherent excitation trapping. These resonances manifest as long-lived ring-down dynamics in the time domain, indicating that the system temporarily stores energy in quasi-stationary Floquet states with suppressed decay, thus prolonging the coherent oscillations observed after the pulse ends.
Figure 13 presents numerical simulations of a driven, interacting four-spin network under a finite-duration pulse. In the experiment, the emitted signal is measured via homodyne detection of the in-phase (I) and quadrature (Q) components, which are proportional to the collective dipole radiation from the system. This output field corresponds to the collective lowering operator , so the measured power reflects the expectation value , directly connecting to the simulated observable . The upper panels show the time-domain evolution of this collective excitation as a function of drive frequency for increasing pulse durations, where pronounced interference patterns emerge in the ring-down following the pulse, as shown in Fig. 13(a). Figure 13(b) displays the corresponding Fourier spectra, which reveal how the spectral content of the ring-down evolves with increasing pulse duration. For short pulses, the abrupt temporal cutoff introduces large energy uncertainty, resulting in broad sidebands. As the pulse duration increases, the drive becomes more spectrally selective, allowing coherent oscillations among near-resonant modes to build up over time. This temporal coherence leads to narrower sidebands in the Fourier spectrum, which progressively sharpen and approach the bare transition frequencies of the system. This sharpening gives rise to the observed V-shaped structure in the frequency domain that we observe experimentally.